# Attosecond-resolved probing of recolliding electron wave packets in liquids and aqueous solutions

### v1

@@ Introduction
Recent experiments have successfully extended HHG to the liquid phase [35, 36, 37, 38, 39, 40, 41], identifying electron recollision as a viable mechanism in bulk liquids [39, 42, 43, 44]. Unlike in gases, where electron quantum paths are dominated by laser parameters, the maximum quantum path length in liquids at ambient pressures is limited primarily by the intrinsic mean free path (MFP) of low-energy electrons and is largely independent of laser wavelength, intensity, or pulse duration [39, 42].
@@
These methods offer two complementary views on the attosecond dynamics of electron dynamics in liquids with 30 as/eV precision.
Remarkably, these two techniques provide independent and matching verifications of the characteristic timescales of ultrafast electron dynamics in solutions, each with its own advantages.
Rather than being coincidental, we argue that this agreement reflects a weak dependence of the ionization times on the emitted photon energy, in agreement with the quantum-orbit description of HHG.
By establishing a mapping of emitted photon energy to the attosecond dynamics of the electronic wavepacket in chemically relevant environments under ambient conditions, this work expands the domain of attosecond HHS into uncharted, but highly relevant regimes.
The ability to track charge motion with sub-cycle precision in liquids lays the foundation for future studies of charge migration in solution, ultrafast solvation dynamics, and light-induced control of liquid-phase chemical reactions, with potentially far-reaching implications for photochemistry, biophysics, and energy science.
@@ Results
We first discuss the two interferometric techniques, starting with Fig. 1a that illustrates active interferometry. This method relies on HHG driven by a phase-controlled two-colour laser field, composed of a fundamental frequency and a weak second-harmonic (2) component. In contrast to a single-colour laser field, where electron quantum paths in consecutive half cycles of the electric field are symmetric and only odd harmonics are emitted [45], the addition of a weak 2 component breaks this symmetry and leads to the generation of even harmonics with an intensity that oscillates periodically with the phase delay between the two colours.
Owing to the different phases accumulated by the electron wave packet in two consecutive half cycles, each even harmonic maximizes when the amount of symmetry breaking along its quantum path is maximal – encoding information about the emission time of the harmonics into the two-colour delay at which the even harmonic maximizes [20, 18].
@@
The schematic of the experimental setup is shown in Fig. 2a. A phase-controlled two-colour (-) laser field with parallel polarizations of the two colours is generated by frequency doubling an 1800-nm fundamental pulse in a -barium borate (BBO) crystal.
The phase delay between the and 2 components is controlled using a pair of fused-silica wedges. The high harmonics generated by this two-colour field in a liquid flat jet or an interchangeable gas jet are detected in a flat-field imaging spectrometer.
@@
The experimental results are further supported by ab-initio time-dependent density-functional-theory (TDDFT) calculations on large water clusters (see methods section for details), which yield a slope of 125 48 as/eV (Fig. 2d).
The magnitude of these slopes is much larger than in the gas-phase measurements, performed under comparable conditions. Due to its higher and longer electron quantum paths, the gas-phase experiment required higher driving intensity ( W/cm2) and a stronger 2 component () to observe a clear modulation of the even harmonics. As shown in Fig. 2c (yellow circles), the extracted slope ( as/eV) is much smaller than in the liquid-phase measurements. This result agrees with calculations based on the SFA for a model system with 10.48 eV, matching that of gas-phase ethanol (orange circles in Fig. 2c). Further experimental and theoretical details are provided in the Methods section.
@@
b TDDFT calculation of the high-harmonic spectrum of a 1-M NaCl solution in water using an 1800-nm driving pulse with a peak intensity of 21013 W/cm2. c Measured yield of H17 (11.7 eV) as a function of NaCl concentration (circles). The orange dashed line is a fit based on a two-species (H2O and Cl-) interference model. d Variation of the relative phase of emission from Cl- and H2O, extracted from a fit to the interference model (left axis) and variation of the electron transit time (right axis), as a function of the emitted photon energy.
In its simplest implementation, this technique requires two species contributing to the harmonic emission. We realize the two-emitter system with aqueous solutions of NaCl, containing H2O, hydrated cations (Na) and anions (Cl), whereby we drop the subscripts “(aq)” for brevity from hereon. Owing to the significantly higher of Na+ (35.4 eV) compared to liquid H2O (11.67 eV) and Cl- (9.55 eV) [52, 53, 54], the cations do not contribute measurably to high-harmonic generation. This statement is confirmed by the experimental observation that the emitted HHG spectra and especially the position of the interference minimum strongly depend on the nature of the anion (Extended Data Fig., EDF 1), but not on that of the cation (EDF 2).
The concentration of Na+ in H2O serves as the control parameter. Note that we use the relative ’s, referenced to the vacuum level, rather than the energy difference between valence bands and the conduction band of the solution, because (i) the ’s are more accurately known than the band gaps and (ii) the relative ’s also reflect the energy difference of the valence bands because of the common conduction band for solute and solvent (see EDF 5d and 6d), in analogy with negatively doped semiconductors [55].
@@
To further demonstrate that this interference reflects a genuine microscopic effect, we performed elaborate ab-initio TDDFT simulations combined with molecular dynamics (MD) of the salt solution (see the Methods section for details), with the results shown in Fig. 3b. The calculations reproduce the observed minimum (upshifted to the next higher order), confirming that it arises from the microscopic electronic response as opposed to macroscopic phase-matching and propagation effects, which are not included in our TDDFT simulations.
While the ab-initio simulations confirm the microscopic origin of the harmonic suppression, performing a fine-grained concentration scan is computationally expensive and time-consuming, and can become intractable for either very low, or very high, solute concentrations (requiring either very large simulation cells, or high level theory for properly capturing the interaction between the solute ions, respectively). For the interpretation, we therefore introduce an analytical two-emitter interference model (sketched in Fig. 3a), which offers an intuitive and transparent way of determining the relative phase shifts and emission-amplitude ratios.
@@
At low concentrations ( M), the validity of the two-emitter model is further supported by our electronic-structure calculations combined with the MD simulations, shown in EDF 5 and 6. They indicate that the HOMO in salt solutions remains highly localized on individual water molecules and anions (EDF 5). Meanwhile, the LUMO is delocalized over the simulation box. The localization of the HOMO ensures that the harmonic emission can be treated as originating from discrete species with well-defined ’s, consistent with the observed interference behavior and supporting the validity of the method chosen for extracting the harmonic phase shift. The delocalization of the LUMO and higher-lying states validates the treatment of the electron dynamics as taking place in a quasi-continuum.
The two-emitter model is also supported by experiments in which we systematically varied the anion. Replacing Cl- with Br- and I- featuring ’s of 9.55 eV, 8.8 eV and 7.9 eV, respectively, and thus an increasing relative to water, we observe that the suppressed harmonic order moves to lower photon energies (EDF 1) – consistent with the fact that lower harmonic orders correspond to the shorter transit times that are necessary to fulfill the destructive-interference condition. Varying the cation has no significant effect, as expected (EDF 2).
@@
b Comparison of the extracted variations of ionization (yellow circles) and emission times (blue circles) with a quantum-orbit calculation (lines) described in the Methods Section, at a peak intensity of 3.31013 W/cm2, for 8.83 eV (bandgap of liquid water). The experimental data has been shifted by constants offsets of 0.7 fs (for ionization time, yellow circles) and 1.9 fs (for the emission time, blue circles), to overlay on the curves of the quantum-orbit calculations. These constant shifts were determined from the mean of the difference of the experimental and quantum-orbit results. The driving wavelength used in all experiments and calculations is 1800 nm.
@@
The attosecond electron dynamics underlying HHG in liquids can be described by three characteristic times: the ionization time (), the transit time (), and the harmonic emission time (). Each of them is encoded in the phases measured in the present work. As explained above, the relative phase of the harmonic emission from H2O and Cl- gives access to the variation of the transit time and the two-colour delays give access to the variation of [20] with photon energy.
Comparing the variations of and in Fig. 4a, we find that they agree within the experimental error. This, in turn, indicates a negligible dependence of the ionization times on the harmonic order. This conclusion is remarkable in that it agrees with the results from gas-phase experiments and the quantum-orbit analysis of gas-phase HHG [16], which both demonstrate that the variation of the ionization time with harmonic order is negligible compared to the variation of the emission time with harmonic order.
@@
Our results have a number of implications. First, the agreement between active and passive interferometry validates both techniques. It reveals that high-harmonic emission from liquids displays a positive linear chirp, which encodes the attosecond dynamics of the underlying electron quantum dynamics. The common energy dependence of both emission and transit times implies a negligible variation of the ionization times with harmonic order, consistent with the quantum-orbit description of HHG.
@@ Methods
MD simulations. To obtain a selection of atomic positions (snapshots) for the TDDFT supercell calculations described below, we performed thermal sampling of the machine learning potential trained by O’Neill and co-workers [64, 65] on DFT forces and energies computed with the dispersion-corrected revPBE-D3 functional [66, 67]. The dynamics were propagated in i-PI [68], using a stochastic velocity rescaling thermostat [69] to equilibrate to a temperature of 300 K. In order to capture nuclear quantum effects in the thermal distribution, we continued sampling using the path-integral molecular dynamics (PIMD) [70] of 12-bead ring polymers coupled to a PIGLET thermostat [71]. The HHG calculations are highly sensitive to the intermolecular structure, and revPBE-D3 slightly overbinds the liquid [72]. To compensate for this effect, we considered PIMD simulations at a selection of densities, namely 0.90, 0.95, 1.00, and 1.10 g/cm3. Overall, HHG spectra derived from snapshots of pure water at 0.90 g/cm3 corresponded best to the experimental signal.
[…]
1.000 g/cm3, and 1.032 g/cm3, respectively, corresponding to experimental densities at 300 K [73], scaled by a factor of 0.9. For every salt concentration, 50 snapshots were recorded at 5 ps intervals and subsequently fed into the HHG calculations.
Ground-state DFT calculations.
We analyzed the electronic density of states (EDOS) and the band structures of the salt solutions, performing ground-state electronic structure calculations on the simulation snapshots using density-functional theory with the PBE [60] and the PBE0 [74] exchange-correlation functionals. The calculations were performed using the FHI-aims software package [75, 76, 77],
with the default tight settings for the basis sets and real-space quadrature grids. The EDOSs were computed at the -point. The total and species-projected EDOSs, obtained with PBE and PBE0 and averaged over all snapshots, are shown in EDF 3. The HOMO and LUMO orbitals, as well as the band structures for selected individual snapshots are shown in EDF 5 (EDF 6) using the PBE0 (PBE) functionals.

### v2

@@ Introduction
Recent experiments have successfully extended HHG to the liquid phase [35, 36, 37, 38, 39, 40, 41], identifying electron recollision as a viable mechanism in bulk liquids [39, 42, 43, 44, 45]. Unlike in gases, where electron quantum paths are dominated by laser parameters, the maximum quantum path length in liquids at ambient pressures is limited primarily by the intrinsic mean free path (MFP) of low-energy electrons and is largely independent of laser wavelength, intensity, or pulse duration [39, 42].
@@
These methods offer two complementary views on the attosecond dynamics of electron dynamics in liquids with 30 as/eV precision. Remarkably, these two techniques provide independent and matching verifications of the characteristic timescales of ultrafast electron dynamics in solutions, each with its own advantages. Rather than being coincidental, we argue that this agreement reflects a weak dependence of the ionization times on the emitted photon energy, in agreement with the quantum-orbit description of HHG. By establishing a mapping of emitted photon energy to the attosecond dynamics of the electronic wavepacket in chemically relevant environments under ambient conditions, this work expands the domain of attosecond HHS into uncharted, but highly relevant regimes. The ability to track charge motion with sub-cycle precision in liquids lays the foundation for future studies of charge migration in solution, ultrafast solvation dynamics, and light-induced control of liquid-phase chemical reactions, with potentially far-reaching implications for photochemistry, biophysics, and energy science.
@@ Results
We first discuss the two interferometric techniques, starting with Fig. 1a that illustrates active interferometry. This method relies on HHG driven by a phase-controlled two-colour laser field, composed of a fundamental frequency and a weak second-harmonic (2) component. In contrast to a single-colour laser field, where electron quantum paths in consecutive half cycles of the electric field are symmetric and only odd harmonics are emitted [46], the addition of a weak 2 component breaks this symmetry and leads to the generation of even harmonics with an intensity that oscillates periodically with the phase delay between the two colours. Owing to the different phases accumulated by the electron wave packet in two consecutive half cycles, each even harmonic maximizes when the amount of symmetry breaking along its quantum path is maximal – encoding information about the emission time of the harmonics into the two-colour delay at which the even harmonic maximizes [20, 18].
@@
The schematic of the experimental setup is shown in Fig. 2a. A phase-controlled two-colour (-) laser field with parallel polarizations of the two colours is generated by frequency doubling an 1800-nm fundamental pulse in a -barium borate (BBO) crystal. The phase delay between the and 2 components is controlled using a pair of fused-silica wedges. The high harmonics generated by this two-colour field in a liquid flat jet or an interchangeable gas jet are detected in a flat-field imaging spectrometer.
@@
The experimental results are further supported by ab-initio time-dependent density-functional-theory (TDDFT) calculations on large water clusters (see methods section for details), which yield a slope of 125 48 as/eV (Fig. 2d). The magnitude of these slopes is much larger than in the gas-phase measurements, performed under comparable conditions. Due to its higher and longer electron quantum paths, the gas-phase experiment required higher driving intensity ( W/cm2) and a stronger 2 component () to observe a clear modulation of the even harmonics. As shown in Fig. 2c (yellow circles), the extracted slope ( as/eV) is much smaller than in the liquid-phase measurements. This result agrees with calculations based on the SFA for a model system with 10.48 eV, matching that of gas-phase ethanol (orange circles in Fig. 2c). Further experimental and theoretical details are provided in the Methods section.
@@
In its simplest implementation, this technique requires two species contributing to the harmonic emission. We realize the two-emitter system with aqueous solutions of NaCl, containing H2O, hydrated cations (Na) and anions (Cl), whereby we drop the subscripts “(aq)” for brevity from hereon. Owing to the significantly higher of Na+ (35.4 eV) compared to liquid H2O (11.67 eV) and Cl- (9.55 eV) [53, 54, 55], the cations do not contribute measurably to high-harmonic generation. This statement is confirmed by the experimental observation that the emitted HHG spectra and especially the position of the interference minimum strongly depend on the nature of the anion (Extended Data Fig., EDF 1), but not on that of the cation (EDF 2). The concentration of Na+ in H2O serves as the control parameter. Note that we use the relative ’s, referenced to the vacuum level, rather than the energy difference between valence bands and the conduction band of the solution, because (i) the ’s are more accurately known than the band gaps and (ii) the relative ’s also reflect the energy difference of the valence bands because of the common conduction band for solute and solvent (see EDF 5d and 6d), in analogy with negatively doped semiconductors [56].
@@
To further demonstrate that this interference reflects a genuine microscopic effect, we performed elaborate ab-initio TDDFT simulations combined with molecular dynamics (MD) of the salt solution (see the Methods section for details), with the results shown in Fig. 3b. The calculations reproduce the observed minimum (upshifted to the next higher order), confirming that it arises from the microscopic electronic response as opposed to macroscopic phase-matching and propagation effects, which are not included in our TDDFT simulations. While the ab-initio simulations confirm the microscopic origin of the harmonic suppression, performing a fine-grained concentration scan is computationally expensive and time-consuming, and can become intractable for either very low, or very high, solute concentrations (requiring either very large simulation cells, or high level theory for properly capturing the interaction between the solute ions, respectively). For the interpretation, we therefore introduce an analytical two-emitter interference model (sketched in Fig. 3a), which offers an intuitive and transparent way of determining the relative phase shifts and emission-amplitude ratios.
@@
At low concentrations ( M), the validity of the two-emitter model is further supported by our electronic-structure calculations combined with the MD simulations, shown in EDF 5 and 6. They indicate that the HOMO in salt solutions remains highly localized on individual water molecules and anions (EDF 5). Meanwhile, the LUMO is delocalized over the simulation box. The localization of the HOMO ensures that the harmonic emission can be treated as originating from discrete species with well-defined ’s, consistent with the observed interference behavior and supporting the validity of the method chosen for extracting the harmonic phase shift. The delocalization of the LUMO and higher-lying states validates the treatment of the electron dynamics as taking place in a quasi-continuum. The two-emitter model is also supported by experiments in which we systematically varied the anion. Replacing Cl- with Br- and I- featuring ’s of 9.55 eV, 8.8 eV and 7.9 eV, respectively, and thus an increasing relative to water, we observe that the suppressed harmonic order moves to lower photon energies (EDF 1) – consistent with the fact that lower harmonic orders correspond to the shorter transit times that are necessary to fulfill the destructive-interference condition. Varying the cation has no significant effect, as expected (EDF 2).
@@
The attosecond electron dynamics underlying HHG in liquids can be described by three characteristic times: the ionization time (), the transit time (), and the harmonic emission time (). Each of them is encoded in the phases measured in the present work. As explained above, the relative phase of the harmonic emission from H2O and Cl- gives access to the variation of the transit time and the two-colour delays give access to the variation of [20] with photon energy. Comparing the variations of and in Fig. 4a, we find that they agree within the experimental error. This, in turn, indicates a negligible dependence of the ionization times on the harmonic order. This conclusion is remarkable in that it agrees with the results from gas-phase experiments and the quantum-orbit analysis of gas-phase HHG [16], which both demonstrate that the variation of the ionization time with harmonic order is negligible compared to the variation of the emission time with harmonic order.
@@
Our results have a number of implications.
First, the agreement between active and passive interferometry validates both techniques. It reveals that high-harmonic emission from liquids displays a positive linear chirp, which encodes the attosecond dynamics of the underlying electron quantum dynamics. The common energy dependence of both emission and transit times implies a negligible variation of the ionization times with harmonic order, consistent with the quantum-orbit description of HHG.
@@ Methods
MD simulations. To obtain a selection of atomic positions (snapshots) for the TDDFT supercell calculations described below, we performed thermal sampling of the machine learning potential trained by O’Neill and co-workers [65, 66] on DFT forces and energies computed with the dispersion-corrected revPBE-D3 functional [67, 68]. The dynamics were propagated in i-PI [69], using a stochastic velocity rescaling thermostat [70] to equilibrate to a temperature of 300 K. In order to capture nuclear quantum effects in the thermal distribution, we continued sampling using the path-integral molecular dynamics (PIMD) [71] of 12-bead ring polymers coupled to a PIGLET thermostat [72]. The HHG calculations are highly sensitive to the intermolecular structure, and revPBE-D3 slightly overbinds the liquid [73]. To compensate for this effect, we considered PIMD simulations at a selection of densities, namely 0.90, 0.95, 1.00, and 1.10 g/cm3. Overall, HHG spectra derived from snapshots of pure water at 0.90 g/cm3 corresponded best to the experimental signal. In light of this, production PIMD simulations of aqueous NaCl solutions at molar concentrations of 0 M (111 molecules in the simulation box), 1 M (97 , 2 NaCl), 3 M (92 , 6 NaCl), and 4 M (89 , 8 NaCl) were performed at the densities of 0.900 g/cm3, 0.933 g/cm3, 1.000 g/cm3, and 1.032 g/cm3, respectively, corresponding to experimental densities at 300 K [74], scaled by a factor of 0.9. For every salt concentration, 50 snapshots were recorded at 5 ps intervals and subsequently fed into the HHG calculations.
Ground-state DFT calculations. We analyzed the electronic density of states (EDOS) and the band structures of the salt solutions, performing ground-state electronic structure calculations on the simulation snapshots using density-functional theory with the PBE [61] and the PBE0 [75] exchange-correlation functionals. The calculations were performed using the FHI-aims software package [76, 77, 78], with the default tight settings for the basis sets and real-space quadrature grids. The EDOSs were computed at the -point. The total and species-projected EDOSs, obtained with PBE and PBE0 and averaged over all snapshots, are shown in EDF 3. The HOMO and LUMO orbitals, as well as the band structures for selected individual snapshots are shown in EDF 5 (EDF 6) using the PBE0 (PBE) functionals.